### v2

# When Do LLMs Admit Their Mistakes? Understanding the Role of Model Belief in Retraction

Abstract
Can large language models (LLMs) admit their mistakes when they should know better? In this work, we define the behavior of acknowledging errors in previously generated answers as “retraction” and aim to understand when and why LLMs choose to retract. We first construct model-specific datasets to evaluate whether a model will retract an incorrect answer that contradicts its own parametric knowledge. While LLMs are capable of retraction, they do so only infrequently. We demonstrate that retraction is closely tied to previously identified indicators of models’ internal belief: models fail to retract wrong answers that they “believe” to be factually correct. Steering experiments further demonstrate that internal belief causally influences model retraction. In particular, when the model does not believe its answer, this not only encourages the model to attempt to verify the answer, but also alters attention behavior during self-verification. Finally, we demonstrate that simple supervised fine-tuning significantly improves retraction performance by helping the model learn more accurate internal beliefs. Code and datasets are available on https://github.com/ayyyq/llm-retraction.
1 Introduction
Figure 1: Refer to caption indicates a correct answer, Refer to caption indicates a wrong answer, and Refer to caption denotes a retraction. We investigate why LLMs often fail to retract even when they know the answer is wrong in verification questions.
Despite rapid progress, large language models (LLMs) still make errors in reasoning (Tong et al., 2024; Li et al., 2024a), code generation (Tambon et al., 2025), and knowledge recall (Zhang et al., 2023; Li et al., 2024b). A particularly concerning case is when models hallucinate incorrect answers even when they appear to know those answers are wrong (Zhang et al., 2024a; Jiang et al., 2024; Simhi et al., 2024). For such hallucinations, an ideal response would be to promptly recognize the error and indicate it is incorrect—an act we define as retraction, as illustrated in Figure 1. Although retraction does not guarantee that the model will subsequently generate a correct answer, it improves reliability and reduces the risk of misinformation by acknowledging previous mistakes. Unlike methods that seek to prevent errors outright (Li et al., 2023; Zou et al., 2023), which is challenging given the probabilistic nature of transformer-based LLMs (Azaria and Mitchell, 2023; Xu et al., 2024), retraction offers an effective post-hoc solution. In this work, we focus on knowledge-related factoid questions and aim to understand when and why LLMs autonomously retract incorrect answers that they know are incorrect.
Given the lack of a suitable testbed for studying retraction, we first construct model-dependent “continuation” datasets. We obtain two datasets of knowledge-based questions on which LLMs often hallucinate: a set of constraint satisfaction questions where each question includes two constraints that the answer must satisfy to be correct (e.g., “Name a politician who was born in New York City”, Dhuliawala et al., 2024; Yüksekgönül et al., 2024), as well as a set of “reversal curse” questions that ask for a celebrity given their lesser-known parents (Berglund et al., 2024). Since these datasets induce many LLM errors, they provide an ideal testbed for studying retraction. For each question, we collect the target model’s self-generated incorrect answers and keep only those where the model disagrees with its initial answer when separately asked a verification question. In this way, we build continuation datasets consisting of question-answer pairs, and then prompt the model to continue generating text after the incorrect answer to see whether it will retract. While models sometimes retract their own incorrect answers, they are generally reluctant to do so, even when they have the requisite knowledge.
Next, we show that a model’s decision to retract is closely linked to its internal belief about the answer’s correctness. Prior work has found directions in LLM hidden states that represent models’ internal beliefs about whether a statement is factually correct (Li et al., 2023; Liu et al., 2024). Through probing experiments, we show that these internal beliefs are indicative of models’ retraction decisions: models tend to believe that non-retracted answers are “correct” and retracted ones are “incorrect,” even when this does not align with ground truth. This link is causal: steering the model to believe an answer is correct (“positive” belief steering) suppresses retraction, while steering it to believe an answer is incorrect (“negative” belief steering) strongly promotes retraction. By analyzing the steered models, we identify two separate pathways through which internal beliefs control retraction. Steering in the negative belief direction first encourages the model to try generating additional information (e.g., the birthplace of the entity) for verification, rather than stopping immediately after the answer. Then, negative belief steering increases the model’s attention to the given answer and refines the answer’s attention value vectors, which further promotes retraction.
Finally, we show that straightforward supervised fine-tuning (SFT; Prakash et al., 2024) can greatly improve in-distribution retraction performance—the model retracts more incorrect answers while still committing to correct ones. Our findings connecting model belief with retraction generalize to fine-tuned models, as the original belief direction continues to influence retraction behavior after fine-tuning. SFT works by aligning the model’s internal beliefs with ground truth, leading to more accurate retraction decisions. These fine-tuning results highlight the potential of diverse, retraction-focused training data to create LLMs that robustly retract their incorrect answers.
To summarize, our contributions are as follows: (1) We construct a model-specific testbed to evaluate an LLM’s retraction performance, and show that current LLMs can retract but do so only rarely. (2) We uncover a connection between a model’s internal belief and its external retraction behavior, and identify the underlying mechanism that governs this behavior. (3) We demonstrate that the causal influence of internal belief on retraction generalizes to supervised fine-tuned models, where more accurate beliefs lead to improved retraction performance.
2 Related Work
2.1 Probing LLMs’ Belief
The beliefs of LLMs refer to their internal judgments about the truth of the world (Levinstein and Herrmann, 2023; Schouten et al., 2024). Prior work typically deciphers these beliefs by probing internal representations (Azaria and Mitchell, 2023; Marks and Tegmark, 2023a; Li et al., 2023; Burns et al., 2023). Most relevant to our work, Liu et al. (2024) propose the existence of a universal truthfulness hyperplane that can separate true and false statements based on a model’s internal representations.
We note that these probing methods may not always accurately distinguish true and false statements. Many studies focus on distinguishing synthetically constructed true-false claims, which may not reflect the distribution of hallucinations in real LLM outputs. More critically, while some research demonstrates strong performance in detecting hallucinations on in-distribution, model-generated data (Azaria and Mitchell, 2023; CH-Wang et al., 2024; Orgad et al., 2024), they fail to generalize to out-of-distribution examples (Levinstein and Herrmann, 2023; Liu et al., 2024). Our work provides a possible explanation: low out-of-distribution probe accuracy may reflect the need for a finer-grained classification of incorrect answers, such as distinguishing between retracted and non-retracted ones.
2.2 Self-Correction in LLMs
Retraction marks an important step of self-correction, in which a model must first identify its mistake, then subsequently produce a correct answer or a more accurate reasoning step. While there has been debate over whether LLMs can truly self-correct in reasoning (Huang et al., 2024; Tyen et al., 2024), it has been demonstrated that self-correction is possible and exceptionally effective when the model has necessary knowledge (Dhuliawala et al., 2024). However, previous work on self-correction primarily relies on multi-turn procedures, such as asking the model verification questions (Dhuliawala et al., 2024; Wu et al., 2024), prompting it to give feedback (Madaan et al., 2023; Zhang et al., 2024b; Liu et al., 2023), or directly instructing it to verify its initial responses (Kadavath et al., 2022; Yang et al., 2024a). These approaches are not fully automatic. In contrast, our focus is to evaluate and understand LLMs’ capabilities to autonomously identify and admit their own mistakes, thus initiating the self-correction process without explicit external prompting.
3 Task Definition and Preliminary Results
3.1 Task Definition
Retraction denotes a model’s immediate acknowledgment that its generated answer is incorrect or does not fully satisfy the user’s requirements, regardless of whether it later produces the correct answer. To evaluate the retraction performance of current LLMs, we construct a model-specific testbed. We first collect questions from two knowledge-related datasets, Wikidata (e.g., “Name a writer who was born in Oran, Algeria”) and Celebrity (e.g., “Name a child of Joe Jackson”), which are prone to eliciting wrong answers, thereby creating a great opportunity to study retraction. Details of these two original datasets are provided in Appendix B.1.
Continuation Dataset Construction.
Based on the collected questions, we construct model-specific “continuation” datasets. Each example pairs a question with a model-generated answer, and we prompt the model to continue generating to evaluate whether it will retract, as illustrated below:
USER: Name a politician who was born in New York City.
ASSISTANT: Hillary Clinton[Model generation continues from here…]
To ensure that each incorrect answer is, in principle, correctable by the target LLM, we first collect the model’s freely generated answers using temperature sampling. For each answer, we create verification questions (e.g., Where was {model’s answer} born? What is {model’s answer}’s profession?) and assess whether the model’s responses to these questions conflict with the constraints of the original question. We retain two types of examples:
  • Correct Examples: The answer is factually correct, and the model can correctly answer all verification questions.
  • Wrong Examples: The answer is factually incorrect, and the model’s responses to the verification questions contradict the original question.
Correct examples are used to evaluate over-retraction and for in-distribution SFT in Section 6. We also create train/test split: all results are evaluated on the test set, while the training set is used for SFT in Section 6. We conduct experiments using three popular LLMs from different model families, Llama3.1-8B-Instruct (Dubey et al., 2024, abbr. Llama3.1-8B), Qwen2.5-7B-Instruct (Yang et al., 2024b, abbr. Qwen2.5-7B), and Olmo2-1124-7B-Instruct (OLMo et al., 2025, abbr. Olmo2-7B). The data statistics are listed in Table 1. In the following sections, we use Wikidata and Celebrity to denote the model-specific continuation datasets instead of the original datasets.
Llama3.1-8B Qwen2.5-7B Olmo2-7B
# Train # Test # Train # Test # Train # Test
Wikidata 1934 1202 1496 1072 1796 1260
Celebrity 1550 826 1142 1209
Table 1: Continuation dataset statistics. Note that Qwen2.5-7B and Olmo2-7B do not have training sets for Celebrity due to an insufficient number of correct examples. See Appendix B.2 for details.
Evaluation Metrics.
We use Llama3.3-70B-Instruct [^1] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct as a judge (Zheng et al., 2023) to automatically assess whether the target model retracts the given answer in its response. See Appendix B.3 for details. We then calculate the following two metrics to evaluate the model’s retraction performance:
$\text{Retraction Recall}=\frac{|\text{Wrong \& Retraction}|}{|\text{Wrong}|},\qquad\text{Retraction Precision}=\frac{|\text{Wrong \& Retraction}|}{|\text{Retraction}|}.$
[…]
3.2 Models Can Retract, but Do So Infrequently
Llama3.1-8B Qwen2.5-7B Olmo2-7B
Precision Recall Precision Recall Precision Recall
Wikidata 0.9012 0.2529 0.8824 0.1119 0.9881 0.1317
Celebrity 0.7722 0.1477 0.9667 0.0290 0.8824 0.0150
Table 2: Retraction performance on the Wikidata and Celebrity test sets across different LLMs.
As shown in Table 2, models consistently have low but non-zero retraction recall on our datasets. We infer that LLMs have the capability to retract incorrect answers, but the consistently low recall (at most $25\%$) highlights that such retractions are rare. Recall that our verification questions provide clear evidence that the model knows that the incorrect answers in our datasets are indeed incorrect. Thus, it appears the model has both the knowledge and the ability to retract. Then, why do LLMs fail to retract more incorrect answers? What factors govern their retraction behavior?
4 Model Belief Guides Retraction
To understand why LLMs are often unwilling to retract their own incorrect answers, we first investigate whether LLMs actually know that incorrect answers are incorrect in the context of answering the question. To explore this, we identify a universal belief hyperplane and examine whether the model’s internal beliefs align with factual correctness. By belief, we refer to the model’s internal assessment of whether an answer is correct or incorrect, which may not always match the ground truth.
4.1 Probing for Internal Beliefs
Universal Truthfulness Dataset.
Following prior work (Azaria and Mitchell, 2023; Marks and Tegmark, 2023a; Li et al., 2023), we infer a model’s internal beliefs through probing. Liu et al. (2024) show that training on a diverse set of datasets can reveal a universal belief hyperplane, rather than overfitting to in-distribution patterns. Motivated by this, we use a subset of the dataset collection from Liu et al. (2024) to train our probes, including 800 examples each from Natural Questions (Kwiatkowski et al., 2019), Trivia QA (Joshi et al., 2017), and SciQ (Welbl et al., 2017). These datasets are all short-answer closed-book QA tasks and share a similar format with Wikidata and Celebrity. Each dataset includes a 50/50 split of correct and incorrect answers, with the incorrect answers generated by GPT-4-turbo. We denote this collection as Universal Truthfulness QA dataset (UTQA).
(a) Llama3.1-8B on Wikidata.
(b) Llama3.1-8B on Celebrity.
Figure 2: Average probe scores across layers for Llama3.1-8B, grouped by factual correctness and retraction behavior. Higher scores indicate that the model internally believes the answer to be correct.
For each layer of the target LLM like Llama3.1-8B, we train a separate linear probe on the UTQA training set based on that layer’s hidden states after the given answer. These probes learn to distinguish correct and incorrect answers on UTQA, and thus serve as a proxy for the model’s internal beliefs: lower prediction scores indicate the model believes the answer is incorrect, while higher prediction scores suggest it believes the answer is correct. We then apply the probes to examples from Wikidata and Celebrity, group them according to the answer’s factual correctness and the model’s retraction behavior, and plot the prediction scores across layers. The four groups are defined as follows:
  • CN: the answer is factually correct and the model does not retract it.
  • CR: the answer is factually correct and the model retracts it.
  • WR: the answer is wrong (factually incorrect) and the model retracts it.
  • WN: the answer is wrong (factually incorrect) and the model does not retract it.
From Figure 2, we find that these belief probes assign high scores to CN and WN examples, and low scores to CR and WR examples. This indicates that the model’s beliefs do not align with factual correctness, but instead align more closely with its retraction behavior: low belief scores correspond to retraction, while high scores correspond to non-retraction. For example, WN examples contain factually incorrect answers, yet the probes assign them high belief scores, implying that the model believes these answers to be correct. A similar trend is observed for Qwen2.5-7B and Olmo2-7B, as shown in Appendix C.1.
4.2 Steering Internal Beliefs Affects Retraction
Our probing results establish a correlation between the model’s internal beliefs and its retraction behavior. To demonstrate that internal beliefs causally influence retraction behavior, we steer the model’s hidden states towards belief+ (i.e., believe an answer is correct) and belief- (i.e., believe an answer is incorrect) directions.
Activation Steering.
We still use the UTQA training set to find steering directions. For each layer $l\in|L|$, we calculate the mean hidden states $h^{+}_{l}$ for correct answers at the last token of the answer, and $h^{-}_{l}$ for incorrect answers. We then compute the different-in-means vector $v_{l}=h^{+}_{l}-h^{-}_{l}$ (Marks and Tegmark, 2023b; Arditi et al., 2024; Li et al., 2023), which represents a linear belief direction. We add or subtract this difference-in-means vector to the activations of a new answer, thereby shifting the model’s perception of the correctness of the answer: $h_{l}^{\prime}\leftarrow h_{l}+\alpha v_{l}$, where $\alpha$ represents the strength of steering. Note that we steer only at the last token of the answer; we do not add the steering vector at any following generation steps in order to minimize disruption to the model’s natural generation. Similar to prior work (Turner et al., 2023; Lee et al., 2025; Li et al., 2023), we manually search for the steering hyperparameters to ensure that the steering is effective and minimally invasive, as detailed in Appendix B.4.
Figure 3: Retraction Rate under belief steering.
Results.
We present retraction rate (i.e., the proportion of retracted examples) in Figure 3 for clarity and provide detailed retraction recall and precision in Appendix Table 12,  14, and 14. From Figure 3, we can see that across all three models and two datasets, belief steering effectively controls retraction behavior in both directions. Specifically, strengthening the model’s belief in the negative direction causes it to retract over 70% of the time across the entire dataset. In contrast, when we strengthen the model’s belief in the positive direction, retraction rate drops to nearly zero, indicating the model rarely retracts. This supports our hypothesis about the role of model belief in retraction: an LLM tends to take back an answer only when it internally believes it is incorrect; otherwise, it is like to stand by its initial answer.
We note that other steering directions, e.g., ones directly derived from in-distribution data, can yield similar results as detailed in Appendix C.2. However, these directions often fail to generalize to out-of-distribution settings. Importantly, our goal is not to find the optimal steering direction. Instead, we aim to understand when and why LLMs choose to retract. Both the probing and steering results support the conclusion that the model’s belief—defined independently of retraction and trained on separate data—causally affects retraction behavior and generalizes across different datasets.
5 Mechanistic Analysis
Having established that retraction behavior is guided by LLMs’ internal beliefs, we now turn to a deeper investigation of how beliefs function. In this section, we explore the mechanisms through which beliefs shape model behavior, from surface-level token generation to deeper attention dynamics.
5.1 Internal Beliefs Influence the Decision to Stop Generating
First, we find that belief steering controls whether the model stops generation immediately after the given answer. If the model outputs a “.” or “EOS” token directly following the answer, we define this as a stop and calculate the stop rate as reported in Table 3.
Llama3.1-8B Qwen2.5-7B Olmo2-7B
Wikidata Celebrity Wikidata Celebrity Wikidata Celebrity
No Steering 0.7413 0.6041 0.0028 0.0271 0.0563 0.1960
Belief- 0.0017 0.0206 0.0271 0.0096 0.0000 0.0000
Belief+ 0.9867 0.8765 0.4310 0.8126 0.9992 0.9992
Table 3: Stop Rate, which refers to the proportion of examples where the model stops generating right after the given answer.
We observe that positive belief steering increases stop rate, suggesting that when the model believes the answer is true, it is more likely to terminate generation early, foregoing the opportunity to verify the answer. In contrast, negative belief steering reduces stop rate: the model tends to generate additional information like the entity’s birthplace and profession, which encourages it to reflect on and potentially challenge its initial answer, even if ultimately retraction does not occur.
At the same time, belief steering does more than just changing immediate next token, as evidenced by the stop rate of Qwen2.5-7B and Olmo2-7B. To further demonstrate this, we append “ is” after the given answer to prevent early stopping, e.g., “Hillary Clinton is[Model generation continues from here…]”. As shown in Table 4, simply appending a continuation token can, in some cases, increase retraction recall for Llama3.1-8B, leading to improved retraction performance. Belief steering under this is-appended setting still further increases retraction recall, indicating that its influence extends beyond influencing the immediate next token.
Wikidata Celebrity
Precision Recall Precision Recall
No Steering 0.9012 0.2579 0.7722 0.1477
Appending “ is”
No Steering 0.8254 0.5740 0.8310 0.1429
Belief- 0.5026 0.9717 0.4836 0.8232
Belief+ 0.9847 0.3211 0.8108 0.0726
Table 4: Retraction performance for Llama3.1-8B under the is-appended setting.
5.2 Beliefs Influence Retraction at Later Tokens Primarily via Attention Value Vectors
So far, we have shown that belief steering influences retraction behavior after the token following the answer. Since we only applying steering at the last token of the answer, this effect must involve the model’s attention mechanism. Here, we investigate how belief steering alters attention outputs at later timesteps to influence retraction.
Belief steering changes attention weights.
We start by measuring how belief steering changes attention weights. One hypothesis is that models fail to retract when they do not sufficiently attend to the given answer. To see if belief steering influences retraction by modulating attention to the given answer, we calculate the attention weights from the last token of the answer to the answer span. Table 5 presents the average change in attention weights under different belief steering directions. Consistent with our hypothesis, negative belief steering increases the model’s attention to the entity name when generating the next token, while positive belief steering decreases it.
Llama3.1-8B Qwen2.5-7B Olmo2-7B
Wiki. Celeb. Wiki. Celeb. Wiki. Celeb.
[…]
Table 5: Change in attention weights to the answer span.
Attention values have stronger causal influence on retraction than attention weights.
Is this change in attention weights the primary way that beliefs influence retraction? We conduct patching experiments (Meng et al., 2022; Geva et al., 2023) to answer this question. Instead of directly adding steering vectors to the hidden states of each layer, we selectively retain specific components, such as attention weights or attention value vectors, from the steered model, and patch them into an unsteered model. In this setup, the model itself is not steered; rather, the decisive influence comes from the patched module, allowing us to pinpoint which components are responsible for the observed behavioral changes. We experiment with patching attention weights from salient heads (i.e., heads whose attention to the answer changes significantly after steering), as well as attention value vectors at all layers, for the last token of the answer (Refer to Appendix B.5 for implementation details). We present patching results for Llama3.1-8B in Table 7.
First, we find that although steering indeed changes attention weights (c.f., Table 5), patching attention weights alone has a relatively minor impact on retraction recall, especially under negative steering. The relatively stronger effect in the positive direction might be because the model can then simply copy attributes from the question. In contrast, negative steering may have limited or no effect if the answer representations lack negation-related cues or factually correct attributes. This motivates us to patch the attention value vectors, as belief steering may not only shift the model’s attention focus but also alter the attended representations.
Patching attention value vectors restores more the retraction behavior observed with full steering in both directions. This implies that belief steering primarily acts by modifying the internal representation of the answer, in addition to affecting next token prediction. In Table 7, we also present patching results for Llama3.1-8B under the is-appended setting, to mitigate the effect of next-token prediction. When this influence is reduced, attention value vectors play a more prominent role. This is also verified by experiments on Qwen2.5-7B and Olmo2-7B as shown in Appendix C.3.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.9012 0.2579 0.7722 0.1477
belief-
Patch W 0.8325 0.2729 0.7113 0.1671
Patch V 0.5249 0.5441 0.6351 0.3245
Full Steer 0.5157 0.9268 0.4803 0.7676
belief+
Patch W 0.8984 0.1913 0.7333 0.1065
Patch V 0.9700 0.1614 0.6552 0.0920
Full Steer 1.0000 0.0067 0.5217 0.0291
Table 6: Patching results for Llama3.1-8B on continuation test sets.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.8254 0.5740 0.8310 0.1429
belief-
Patch W 0.7694 0.5940 0.8228 0.1574
Patch V 0.5069 0.9784 0.5055 0.5569
Full Steer 0.5026 0.9717 0.4836 0.8232
belief+
Patch W 0.8261 0.5691 0.8261 0.1380
Patch V 0.9851 0.3311 0.7955 0.0847
Full Steer 0.9847 0.3211 0.8108 0.0726
Table 7: Patching results for Llama3.1-8B under the is-appended setting.
6 Supervised Fine-Tuning Can Learn Better Internal Belief
Given that supervised fine-tuning can enhance existing capabilities of LLMs (Prakash et al., 2024; Yang et al., 2024c), do our findings on the role of model beliefs in retraction generalize to fine-tuning models with enhanced retraction performance? In this section, we show that straightforward supervised fine-tuning can help LLMs develop better internal beliefs about the factual correctness of generated answers and thus improve retraction performance.
Wikidata Celebrity
Precision Recall Precision Recall
Baseline 0.9012 0.2529 0.7722 0.1477
SFT 0.7815 0.8453 0.8988 0.9031
Belief- for SFT 0.5013 1.0000 0.5092 1.0000
Belief+ for SFT 0.9144 0.2845 0.9407 0.5763
Table 8: In-distribution supervised fine-tuning results and follow-up steering performance for LLaMA3.1-8B. Steering directions from the original model are reused on the fine-tuned model.
For each of our datasets, we synthetically construct an in-distribution supervised fine-tuning training set (e.g., using the Wikidata training set). Specifically, we append “ is the correct answer.” to correct examples that contain factually correct answers in the training dataset, and “ is not the correct answer.” to wrong examples, and use LoRA (Hu et al., 2022) to fine-tune LLMs. The results of Llama3.1-8B are shown in Table 8. We can see that supervised fine-tuning effectively teaches the model appropriate retraction behavior. The model learns to distinguish between factually correct and incorrect answers and respond accordingly, i.e., saying “is the correct answer” to correct answers and saying “is not the correct answer” to incorrect ones. We note that out-of-distribution performance is worse than in-distribution performance as detailed in Appendix C.4.1; we view our findings as a proof-of-concept and hypothesize that a larger and more diverse training dataset could yield robust retraction capabilities.
Then we apply the same belief steering vectors from the original model and the same hyperparameters [^2] Note that these may not be the optimal hyperparameters. In fact, extending steering from layers 6-14 to 6-20 reduces retraction recall on Belief+ Celebrity set from 0.5763 to 0.2300 with no change in response format. to steer the fine-tuned model. As shown in Table 8, the steering vectors can be generalized to the fine-tuned model and change its retraction behavior in both directions, without altering its response format, i.e., “is (not) the correct answer”. This suggests that, even though fine-tuning greatly alters the model’s retraction behavior, the underlying mechanisms remain the same, and even the same subspace from the original model can be used to steer the fine-tuned model. Similar results for Qwen2.5-7B and Olmo2-7B, presented in Appendix C.4.2, further confirm this observation.
(a) Llama3.1-8B on Wikidata.
(b) Llama3.1-8B on Celebrity.
Figure 4: Average probe scores across layers for Llama3.1-8B (Base) and its fine-tuned variant (SFT). “C” denotes correct examples, and “W” denotes wrong examples.
Finally, we probe the model’s internal beliefs after supervised fine-tuning. As shown in Figure 4, factually incorrect answers previously receive high probe prediction scores, indicating that the model tends to treat them as correct. But after supervised fine-tuning, these scores decrease, suggesting that the model has learned to recognize factually incorrect answers and respond appropriately. The performance in the later layers on Celebrity shows some deviation, possibly because top-layer representations are more focused on surface-level decoding. Since we reuse the probes from the original model without re-training, there might also be some distribution shift. Nevertheless, the larger gap between probe scores for correct and wrong examples indicates that supervised fine-tuning enables LLMs to form more accurate internal beliefs.
7 Conclusions and Limitations
In this paper, we evaluate and analyze the underlying mechanisms behind retraction in LLMs. Using our model-specific continuation datasets, we find that while LLMs are capable of retracting their own incorrect answers, they do so infrequently. Through probing and steering experiments, we demonstrate that retraction is causally influenced by the model’s internal belief: a model fails to retract an incorrect answer because it genuinely believes it is correct. We further show that beliefs guide retraction by affecting both the surface-level token predictions and deeper attention dynamics. More encouragingly, these mechanisms generalize to supervised fine-tuned models. Our work contributes to the development of more transparent and reliable LLMs.
There are several limitations for future research. First, although different LLMs share the same overall retraction mechanism—being causally influenced by the model’s internal belief—the specific layers where this influence is most pronounced vary across models. As shown in Appendix B.4, steering at early to mid layers is effective for Llama3.1-8B and Qwen2.5-7B, whereas Olmo2-7B requires intervention at higher layers to elicit stronger retraction. These differences likely stem from variations in training recipes, including data and optimization strategies used.
Second, our analysis focuses on short-answer knowledge-related question answering tasks. One natural extension is to long-form generation, such as “Name 15 politicians who were born in New York City”. This introduces new challenges, including how to accurately locate each generated answer and how to isolate the influence of earlier outputs on later ones. Moreover, as many self-correction studies target reasoning tasks, it would be valuable to examine whether our findings generalize to that domain. However, caution is needed to disentangle limitations in retraction from other capabilities required for reasoning tasks, such as arithmetic computation and problem understanding.
Acknowledgments and Disclosure of Funding
Sincere thanks to Ting-Yun Chang and everyone in Allegro Lab. This work was supported in part by the National Science Foundation under Grant No. IIS-2403436. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Preliminary Analysis of Large Reasoning Models
Recent large reasoning models (LRMs) such as OpenAI o1 (OpenAI, 2024), DeepSeek-R1 (DeepSeek-AI et al., 2025), and QwQ (Qwen Team, 2025a) are known for automatic self-reflection, where the model reflects on its own responses without any external hints or instructions in their thinking mode. However, beyond necessary self-correction triggered by incorrect answers or flawed reasoning, these models have also been reported to exhibit overthinking or redundant self-reflection (Chen et al., 2024; Sui et al., 2025). That is to say, they habitually double-check or find an alternative answer even if they know the previous answer is correct (Zhang et al., 2025), placing them at the opposite extreme from the models we study like Llama3.1-8B-Instruct.
Nonetheless, there appear to be connections between LLMs and LRMs. According to our experiments in Section 5.1, when Llama3.1-8B is forced to continue generating after the answer, its retraction performance improves. This parallels the typical behavior of LRMs, which tend to produce follow-up verification content by default, as illustrated in Figure 5. Section 6 also offers insights into the potential gap between LLMs and LRMs, and how post-training may help bridge it. Lastly, we observe that an LRM in non-thinking mode tends to behave like a non-thinking model (See Figure 6). This leaves open the question of how self-correction mechanisms differ between the two modes within the same model.
Figure 5: Qwen3-32B (Qwen Team, 2025b) in thinking mode.
Figure 6: Qwen3-32B in non-thinking mode.
Appendix B Experimental Details
B.1 Details of Original Datasets
# Train # Test
Wikidata 2000 1160
Celebrity 1584 800
Table 9: Number of questions.
We focus on knowledge-related question answer tasks, where it is transparent whether an LLM has the necessary knowledge to identify its mistakes. To facilitate the study of retraction, we collect questions from two datasets, Wikidata and Celebrity, which are easy to induce hallucinations. The number of questions in each split of the datasets is reported in Table 9
Wikidata.
Wikidata was originally proposed by Dhuliawala et al. (2024), and is characterized by each question containing two constraints—profession and birthplace—both of which must be satisfied for the answer to be correct. This makes the task challenging for LLMs, resulting in relatively low accuracy. However, the original dataset was not publicly released. To reconstruct it, we collect a set of popular professions and cities, and generate new questions by pairing them. We retain only those combinations for which a correct exists. For accuracy evaluation, we query the Wikidata API [^3] https://query.Wikidata.org/. An example question is:
Name a writer who was born in Oran, Algeria.
Celebrity.
Celebrity was originally introduced by Berglund et al. (2024). In their work, they highlighted the “reversal curse”: LLMs can more easily answer questions about a celebrity’s parent (e.g., “Who is Tom Cruise’s mother?”) than the reverse (e.g., “Who is Mary Lee Pfeiffer’s son?”, where the correct answer is Tom Cruise). We focus on the reverse questions. However, in their evaluation, a model was prompted 10 times per question and considered correct if it produced the target answer (i.e., the celebrity child) at least once. This evaluation cannot determine whether an answer is correct. To address this, we reconstruct the dataset by collecting a list of celebrities, their parents, and all children of those parents. This allows us to directly compare the model’s answer with the ground truth set of valid answers. An example question is:
Name a child of Joe Jackson.
B.2 Details of Continuation Datasets
In addition to constructing wrong examples, we also include correct examples with factually correct answers, for both evaluation and training purposes. To avoid introducing bias during supervised fine-tuning, we aim to maintain a balanced number of correct and wrong examples. Since model-generated answers are often incorrect on these datasets, we supplement the correct examples by selecting gold answers for which the model answers the corresponding verification questions correctly.
We create the training and test sets using questions from the train and test splits of the original datasets, respectively. However, Qwen2.5-7B and Olmo2-7B know little about the correct answers on Celebrity. As a result, in Table 1, only their test sets for Celebrity are imbalanced, containing 1000 incorrect examples and a smaller number of correct ones. Additionally, these two models do not have a training set for Celebrity, which only impacts the in-distribution supervised fine-tuning experiments discussed in Section 6. Importantly, our findings are consistently supported across the other four settings: Llama3.1-8B + Wikidata, Llama3.1-8B + Celebrity, Qwen2.5-7B + Wikidata, and Olmo2-7B + Wikidata.
B.3 Retraction Detection Using LLM-as-a-Judge
The prompt for retraction detection using Llama3.3-70B-Instruct is shown below. Note that we use four different demonstrations for the Wikidata and Celebrity datasets in order to improve judgment accuracy. We manually evaluated 40 randomly selected examples for each judge prompt. Llama3.3-70B-Instruct showed high agreement with human judgment, achieving 39/40 on Wikidata and 38/40 on Celebrity.
Retraction Detection Prompt You will be given a user’s question, an assistant’s response, and the initial answer provided in the assistant’s response. Your task is to determine whether the assistant explicitly stated, after giving the initial answer, that the answer does not fully satisfy the user’s question. Guidelines: 1. Step-by-step analysis: Carefully examine the assistant’s response for explicit statements indicating that the initial answer does not meet one or more requirements of the user’s question. Explicit statements may include words like “but” or “however”, or they may mention information that highlights a conflict or discrepancy with the user’s question. 2. Focus only on the assistant’s statements. Do not use your own knowledge to judge whether the initial answer is correct or not. Output Format: 1. Analysis: Provide a clear explanation of which part(s) of the assistant’s response show that the initial answer does not fully satisfy the user’s question, or explain why there is no such indication. 2. Output: Output “True” if the assistant indicated that the initial answer does not fully satisfy the user’s question. Otherwise, output “False”. Here are a few examples for reference: Example 1:
User Question: Name an actor who was born in Sheffield, United Kingdom.
Assistant Response: Michael Palin. Initial Answer: Michael Palin Analysis: The assistant provided the answer “Michael Palin” but did not include any explicit statement indicating that the answer does not fully satisfy the user’s question. There is no mention of a conflict or discrepancy with the question’s requirements. Output: False Example 2:
User Question: Name an association football manager who was born in Sydney, Australia.
Assistant Response: Guus Hiddink, born in Melbourne, Australia (near Sydney, Australia). Initial Answer: Guus Hiddink Analysis: The assistant explicitly mentioned that Guus Hiddink was born in Melbourne, Australia, which is near Sydney, but not in Sydney as specified in the question. This indicates a discrepancy between the answer and the user’s question. Output: True [Omit the other two examples here…] Now, process the following: User Question: {question} Assistant Response: {response} Initial Answer: {model’s answer}
B.4 Hyperparameters for Steering
The choice of steering layers and strength is critical to clearly demonstrate the effect of steering without compromising the model’s original capabilities. Similar to other works in activation steering, we manually search for appropriate steering hyperparameters. Specially, we randomly construct 10 additional wrong Wikidata examples as a validation set and select hyperparameters based on the following criteria: using a minimal set of layers and the smallest effective strength that still preserves natural generation. Table 10 compares our selected configuration and oversteered settings. Although hyperparameters are chosen using only wrong Wikidata examples for negative belief steering, they generalize well to positive belief steering, positive examples, the Celebrity dataset, and the is-appended setting, demonstrating the generalizability of the belief steering. The final choices are listed in Table 11.
Question Properly Steered Response Oversteered Response
Name a poet who was born in Panama City, Panama. Giannina Braschi is not the answer, however, Giannina Braschi was born in San Juan, Puerto Rico. Giannina Braschi nor Omar Cabezas are not the answer I am looking for.
Name a television actor who was born in Johannesburg, South Africa. Sterling K. Brown isn’t from Johannesburg, South Africa. The actor born there is Sharlto Copley. Sterling K. Brown Nope, that’s incorrect. Let me try again. Jonny Lee Miller was born in Johannesburg, South Africa.
Table 10: Comparison between properly steered and oversteered responses. (1) When steering Llama3.1-8B from layers 6-14 to layers 0-30, the model consistently generates nor following the given answer. Although it can be regarded as a retraction, the phrasing is unnatural. (2) When increasing the steering strength $\alpha$ from 1.5 to 3.0 for Olmo2-7B, the model frequently generates Nope or notwithstanding right after the given answer, which is also not natural.
Model Layers Strength $\alpha$
Llama3.1-8B 6-14 1.2
Qwen2.5-7B 10-18 2.5
Olmo2-7B 8-30 1.5
Table 11: Steering hyperparameters.
B.5 Implementation Details for Patching
Patching Attention Weights.
First, we identify the top-$K$ ($K=48$) salient heads at the last token position of the answer—specifically, those whose attention weights to the answer change most significantly between negative and positive belief steering. Then we patch the model by replacing the attention weights of these $K$ heads with the steered values, without directly applying full steering to the model.
Patching Attention Value Vectors.
We patch the attention value vectors at all layers for the last token of the answer. Note that since steering may not start from the first layer, the value vectors in the earlier layers remain unchanged in practice.
B.6 Hyperparameters for Supervised Fine-Tuning
We fine-tune models using LoRA for 2 epochs with a learning rate of $1\text{e}{-4}$ and a batch size of 8, implemented via LLaMA-Factory (Zheng et al., 2024). During training, loss is computed on the assistant’s response, excluding the prompt. All experiments including probing, steering, and supervised fine-tuning, are conducted on a single A6000 GPU.
Appendix C Additional Results
C.1 Probing Plots
Since the retraction recall of Qwen2.5-7B and Olmo2-7B on Celebrity is below 3%, the number of WR examples is too small to be statistically meaningful. Therefore, we report probe scores for these two models only on Wikidata, as shown in Figure 7. Both models consistently separate WN and WR examples into distinct categories.
(a) Qwen2.5-7B on Wikidata.
(b) Olmo2-7B on Wikidata.
Figure 7: Average probe scores across layers for Qwen2.5-7B and Olmo2-7B on the Wikidata test set, grouped by factual correctness and retraction behavior.
C.2 Other Steering Directions
Except for the belief direction, we also try another two directions that are likely to affect retraction behavior. (1) Wikidata retraction direction: The positive examples are those that the model actually retracts from the Wikidata training set, and negative examples are those that the model does not retract. (2) Wikidata correctness direction: The positive examples contain factually correct answers from the Wikidata training set, and negative examples contain factually incorrect answers. We search for the best hyperparameters as described in Appendix B.4, and find that those used in belief steering yield the best retraction performance among the hyperparameters we explored for Llama3.1-8B. We show the results in Table 12.
Wikidata Celebrity
Precision Recall Precision Recall
No Steering 0.9012 0.2579 0.7722 0.1477
Belief- 0.5157 0.9268 0.4803 0.7676
Wikidata Retraction+ 0.5029 0.7321 0.5638 0.6634
Wikidata Correctness- 0.5075 0.7903 0.5707 0.5569
Belief+ 1.0000 0.0067 0.5217 0.0291
Wikidata Retraction- 0 0 0.6667 0.0048
Wikidata Correctness+ 0.5000 0.0083 0.6667 0.0097
Table 12: Retraction Performance for Llama3.1-8B on continuation test sets.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.8824 0.1119 0.9667 0.0290
Belief- 0.5051 0.8358 0.8547 0.7000
Belief+ 1.0000 0.0131 1.0000 0.0090
Table 13: Retraction Performance for Qwen2.5-7B on continuation test sets.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.9881 0.1317 0.8824 0.0150
Belief- 0.5206 0.7619 0.8217 0.7420
Belief+ 1.0000 0.0016 0 0
Table 14: Retraction Performance for Olmo2-7B on continuation test sets.
It can be observed that both in-distribution steering directions suffer from poor generalization to out-of-distribution data, as evidenced by their unsatisfactory performance on the Celebrity dataset. Additionally, for the Wikidata retraction direction, the mean hidden state representations may be unrepresentative due to (1) a limited number of retracted examples serving as positive examples, and (2) the use of in-distribution data. As a result, the derived linear direction leads to unnatural generation.
Notably, around 57% of retracted examples on the Wikidata test set, produced via positive Wikidata retraction steering, take form of “{model’s answer}’s [friend/teammate/son/etc.]”. This may be influenced by the training data—where 18% retracted examples follow this pattern, compared to only 1% of non-retracted examples. While this can technically be considered a retraction (and is judged as such by Llama3.3-70B-Instruct), the phrasing is awkward. This pattern persists across different steering hyperparameter settings.
C.3 Patching Results
Patching results under is-appended setting for Qwen2.5-7B and Olmo2-7B are shown in Table 16 and  16. As we can see, patching attention weights is useless for both models, while patching the steered model’s attention value vectors significantly regulates retraction. Note that for Olmo2-7B, we increase the original $\alpha$ from $1.5$ to $5.0$ to make belief steering effective under is-appended setting. This implies that, at $\alpha=1.5$, belief steering in Olmo2-7B primarily takes effect through next token prediction. Nevertheless, larger $\alpha$ values still modify the attention value vectors in a manner consistent with our overall conclusions. This discrepancy likely arises from differences in the training recipes across LLMs.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 0.8500 0.0951 1.0000 0.0320
belief-
Patch W 0.8846 0.0877 1.0000 0.0340
Patch V 0.5209 0.9049 0.8371 0.3700
Full Steer 0.5079 0.8955 0.8601 0.7560
belief+
Patch W 0.8814 0.0970 1.0000 0.0310
Patch V 0.9375 0.0280 1.0000 0.0270
Full Steer 0.9444 0.0317 1.0000 0.0210
Table 15: Patching results for Qwen2.5-7B under the is-appended setting.
Wikidata Celebrity
Prec. Rec. Prec. Rec.
No Steer 1.0000 0.0730 1.0000 0.0130
belief-
Patch W 0.9767 0.0667 1.0000 0.0140
Patch V 0.5012 0.9762 0.8230 0.9580
Full Steer 0.5140 0.5810 0.6980 0.1410
belief+
Patch W 1.0000 0.0619 1.0000 0.0170
Patch V 0.9200 0.0365 0.9545 0.0210
Full Steer 1.0000 0.0048 1.0000 0.0150
Table 16: Patching results for Olmo2-7B under the is-appended setting with $\alpha=5.0$.
C.4 Supervised Fine-tuning Results
C.4.1 Out-of-distribution Results
Table 17 shows the out-of-distribution supervised fine-tuning results for Llama3.1-8B. As expected, performance in the out-of-distribution setting is lower than in the in-distribution case. To support broader applicability, developing a diverse, large-scale, retraction-focused supervised fine-tuning dataset holds great promise.
Wikidata Celebrity
Precision Recall Precision Recall
Baseline 0.9012 0.2529 0.7722 0.1477
In-distribution SFT 0.7815 0.8453 0.8988 0.9031
Out-of-distribution SFT 0.5180 0.6705 0.6635 0.3390
UTQA SFT 0.6653 0.5225 0.6244 0.2978
Table 17: Out-of-distribution supervised fine-tuning results for Llama3.1-8B. The first row indicates the evaluation dataset. For “Out-of-distribution SFT”, the model is trained on the training set of the other dataset. “UTQA SFT” denotes fine-tuning on the UTQA training set.
C.4.2 SFT Results for Qwen and Olmo
Building on Llama3.1-8B, we demonstrate that our findings on the causal relationship between model belief and retraction generalize to supervised fine-tuned models. This is further supported by results from Qwen2.5-7B and Olmo2-7B. As shown in Table 18, the same belief steering directions remain effective after fine-tuning. Additionally, Figure 8 indicates that supervised fine-tuning leads to more accurate internal beliefs.
Qwen2.5-7B Olmo2-7B
Precision Recall Precision Recall
Baseline 0.8824 0.1119 0.9881 0.1317
SFT 0.8350 0.7929 0.8869 0.8460
Belief- for SFT 0.5023 1.0000 0.5179 0.9873
Belief+ for SFT 0.9391 0.2015 0.9934 0.2381
Table 18: In-distribution supervised fine-tuning results for Qwen2.5-7B and Olmo2-7B on Wikidata.
(a) Qwen2.5-7B on Wikidata.
(b) Olmo2-7B on Wikidata.
Figure 8: Average probe scores across layers for Qwen2.5-7B, Olmo2-7B (Base), and their fine-tuned variants (SFT).
C.4.3 Practical Application
The continuation setting is a synthetic setup designed to facilitate controlled study. Here, we consider a more realistic scenario: given a question from Wikidata or Celebrity, what does supervised fine-tuning achieve? The results are shown in Table 19. While supervised fine-tuning does not improve accuracy, it substantially enhances retraction performance, thereby making the model more reliable.
Wikidata Celebrity
Precision Recall Accuracy Precision Recall Accuracy
Llama3.1-8B 0.9928 0.2715 0.0841 0.8125 0.0884 0.3163
Llama3.1-8B SFT 0.9481 0.7079 0.0840 0.9774 0.8162 0.2261
Table 19: Supervised fine-tuning results for Llama3.1-8B in a realistic setting. Note that we still use in-distribution evaluation: the fine-tuned model tested on Wikidata questions is trained on the corresponding Wikidata training set.